Operations Report – January 21st

Crew 203 Operations Report 21-JAN-2019

SOL: 9

Name of person filing report: Freddy Castaneda

Non-nominal systems: Nothing new to report.

Notes on non-nominal systems: None

Generator (hours run): 5hr 00min; Turned on last night (20Jan2019) at 17:40; Turned off about 20:40; It was shuted down all night long; This morning (21Jan2019) it was Turned On from 9:00 to 11:00, it was running nominally. At 14:00 the SOC was almost 60%, it was a pretty cloudy day with some snow shower, so the Generator was turned On. It ran for 2 hours and the SOC display was 100%. It wasTurned On again tonight (21Jan2019) at 17:40 and is running nominally.

Solar SOC – Turned on (20Jan2019) 85%; Turned off (21Jan2019) 50%; Turned on (21Jan2019) 86%

Diesel Reading – Still going under 1/2 tank.

Propane Reading – Almost 20%

Ethanol Free Gasoline – Nothing to report.

Water (auxiliary tank) – Not in use.

Water (static tank) – About 40%; 215 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – Over 49%; 162 gallons

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 18:30 to refill tank

Water Meter: 01403355

Toilet tank emptied: No.

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Used for EVA to the Moons.

Hours: 69.6

Beginning charge: 100%

Ending charge: 38%

Currently charging: Yes

Opportunity rover used: Used to test brakes.

Hours: 46.0

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Used for EVA to the Moons.

Hours: 71.5

Beginning charge: 100%

Ending charge: 68%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly.

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used.

CrewCar used and why, where? Off campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Suit #3 still pending for review. Suit #11 quit during EVA. Currently charging.

Summary of Hab operations: Today we realized that half of the toilet paper provided is actually kitchen towels. 🙁

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing to report.

EVA Report – January 21st

Crew 203 EVA Report 21-01-2019

EVA # 5

Author: Liza Forero

Purpose of EVA:

Geologic recognition of the zone. Take samples and data for microbiology test and take some physic – chemical parameters of the ground using a LabQuest. The main parameters that are going to be taken are conductivity, pH and temperature.

Geology recognition of outcrops, topographic highs, sediment, canyons and lithology.

Start time:
12:11

End time:
16:00

Narrative:
 The EVA was successful, despite the weather, that was very windy and for a couple of minutes that rain, our EVA could be develop in the best possible way. We arrived at White Moon, leave the rovers over the Cow Dung Road and started to walk at west over the Brahe Highway. During the walk we covered some topographic highs. Finally, we arrived at the Beige Moon where we took some data for the physic – chemical parameters. 
During the walk we took samples that are going to be analyzed at the laboratory with the LabQuest. We described the lithology or the zone and found different types and colors of clays. 
During the EVA there were no trouble with rovers but at the middle of the simulation the suit of our commander Oscar Ojeda stopped working so he had to break simulation, but the rest of us could continue with the activities in a nominal way. After, we arrived at the Beige Moon and finish the activity we went back to the rovers and go back to the HAB. 
During the walk we were taking pictures and the register of the stratigraphy of the zone.

While we were walking we felt how the clays were so humid and plastic because of the water they absorbed with the snow and melted water.

Destination: Beige Moon (vehicle and walking).

Coordinates: 516500 E, 4254290 N

Participants:
Liza Forero, Yael Méndez, Oscar Ojeda, Hermes Bolivar

Road(s) and routes per MDRS Map: Cow Dungroad to the north till we get to the white moon, after that we are going to take west in the Brahe Highway to arrive to the Beige Moon.

Mode of travel: Walking and Driving

Vehicles used: Spirit and Curiosity

Greenhab Report – January 21st

Crew 203 Greenhab Report 21-Jan-2019
Greenhab Officer:
Hermes Bolivar

Environmental Control:
Heating
Cooling w/ ambient air (0 hrs)

40% Shade Cloth on
80% Shade Cloth on

Average Temperatures:
Low: 16.6°C
High: 27.1°C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
16 gallons

Water in Blue Tank (gallons):
55%, 165 Gallons.

Times of Watering for Crops:

11:20
14:00
17:12
19:16

Changes to crops:

* The spinach has begun to emerge.

Narrative:

I Harvest carrot, however, some plants need to grow more, i will harvest its the next days.

Harvest:

* 66 g of carrot.
*2 g of cilantro.

Support/Supplies Needed:

NONE

Journalist Report – January 21st

Sol 9

Title: Game Over, Insert Coin

Author: Yael Natalia Méndez / Crew Scientist and HSO

A night of total darkness, full of fog and stories of terror, accompanied our dreams. It is curious that all the crew have had surreal dreams where they combine the reality that awaits us outside with the daily life of the station. It dawned snowing, it doesn’t stop being worrying the uncertainty of knowing if we can make the EVA, not only for the science that we want to do but for the opportunity to go out to "breathe fresh air".

With a finished puzzle and the missing Table topics cards, we started the routine. We go a little more than half of the rotation and fatigue is evident, the spirits a little low, however, we are all doing our best to continue making the most of this experience. Today I would like to embrace my family and my friends, I have read their messages and at this moment they can not imagine how important it is to receive them, that little greeting renews the spirit and gives strength to continue.

With an approved EVA, we started the journey to the moons. With the weather conditions against us, we raffle the situation well, we have great samples. We marvel at landscapes of colored clays, resembling the vegetation of our beautiful Colombian mountains. The cold did not stop us, our bodies were tired, frozen, however, none wanted to return yet, we all wanted to see that immensity. Once back home, what we now call home, we only hoped to have the warmth of a "aguapanelita". Only 9 soles have passed and we already want to eat "arepa", "mango biche", "bom bom bum", even though we have enjoyed delicious Colombian dishes like the beans of yesterday and the “calentao” of this morning.

Those things are undoubtedly important, but what made me reflect today was the fact that Mars is a hostile environment to preserve human life, how would we make the right decisions to survive? If the life of a friend were at risk, how would we act? Today I’m really happy because our EVA is part of a simulation and we all return home …

Sol 9

Title: Game Over, Insert Coin

Autor: Yael Natalia Méndez / Científica y Oficial de Salud y Seguridad.

Una noche de total oscuridad, llena de bruma e historias de terror, acompañó nuestros sueños. Es curioso que todos los tripulantes han tenido sueños surrealistas donde combinan la realidad que nos aguarda afuera con la cotidianidad de la estación. Amaneció nevando, no deja de ser preocupante la incertidumbre de saber si podemos realizar la EVA, no solamente por la ciencia que queremos hacer sino por la oportunidad de salir a “respirar aire fresco”.

Con un rompecabezas terminado y las cartas de Table topics desaparecidas, empezamos la rutina. Vamos un poco más de la mitad de la rotación y el cansancio es evidente, los ánimos un poco bajos, sin embargo, todos estamos haciendo nuestro mejor esfuerzo para seguir aprovechando esta experiencia al máximo. Hoy quisiera abrazar a mi familia y a mis amigos, he leído sus mensajes y en este momento no pueden imaginar que tan importante es recibir estos, ese pequeño saludo renueva el espíritu y da fuerzas para continuar.

Con una EVA aprobada, comenzamos la travesía hacía las lunas. Con las condiciones climáticas en nuestra contra, sorteamos bien la situación, tenemos magnificas muestras. Nos maravillamos con paisajes de arcillas de colores, asemejando la vegetación de nuestras bellas montañas colombianas. El frío no nos detuvo, nuestros cuerpos estaban cansados, congelados, sin embargo, ninguno quería regresar aún, todos queríamos divisar esa inmensidad. Ya de regreso a casa, la que ahora llamamos casa, solo esperábamos tener la calidez de una “aguapanelita”. Solo han pasado 9 soles y ya queremos comer “arepa”, “mango biche”, “bom bom bum”, a pesar de que nos hemos deleitado con ricos platos colombianos como la frijolada de ayer y el “calentao” de esta mañana.

Esas cosas son importantes indudablemente, pero lo que hoy me hizo reflexionar fue el hecho de que Marte tiene un ambiente hostil para preservar la vida humana, ¿cómo tomaríamos las decisiones adecuadas para sobrevivir? ¿Si la vida de un amigo estuviera en riesgo, como actuaríamos? Hoy realmente estoy feliz porque nuestra EVA es parte de una simulación y regresamos todos a casa.

Sol Summary – January 21st

Crew 203 Sol Summary Report 21-Jan-2019

Sol: 9

Summary Title: Aeolus Clash

Author’s name: Oscar Ojeda

Mission Status: Nominal

Sol Activity Summary: We woke up to check the SOC of the generator, given it’s anomalous functioning. We had a Colombian breakfast, and then worked for a while in our respective projects, before getting ready for the EVA. It was very windy outside, and during the EVA one of the suits shut off. We returned to the hab had lunch and rested a little, before starting to work on the reports. We expect to have a hot dinner to help with the cold night.

Look Ahead Plan: We’ll keep the EVAs going, and processing the results of the science projects.

Anomalies in work: The diesel generator was turned off during the night. During the EVA, one of the suits shut down.

Weather: WINDY, very windy, cloudy (nimbus and cumulonimbus) and with sporadic rain.

Crew Physical Status: In good health.

EVA: 1, a team of 4 went to the beige moon region.

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, Journalist Report, EVA Request, EVA Report.

Support Requested: None